OpenStudy (anonymous):

write and equation for the line through (-5,2) with a slope of 0

OpenStudy (anonymous):

y - 2 = 0(x - (-5)) => y = 2.

OpenStudy (anonymous):

The answer I get y=-5

OpenStudy (anonymous):

y = - 5 is incorrect. See my reply.

OpenStudy (anonymous):

My mistake! You are correct! The answer is Y=2.

OpenStudy (anonymous):

yeah, if you're given a slope and a point just plug into point slope formula. \[y-y_1 = m(x-x_1) \text{ where } x_1,y_1\] is the x,y values of the point and m is the slope